What is the [H3O-] and [OH-] ion concentration respectively in a solution that contains 0.155 M HNO3?



1
Expert's answer
2021-03-01T03:25:19-0500

Solution:

HNO3 + H2O → H3O+ + NO3-

After complete dissociation: CM(HNO3) = [H3O+] = [NO3-]

So concentration of [H3O+] = 0.155 M

[H3O+] = 0.155 M


For any aqueous solution at 25C:

Kw = [H3O+] × [OH-] = 1×10-14

Hence,

[OH-] = Kw / [H3O+]

[OH-] = (1×10-14) / (0.155) = 6.45×10-14

[OH-] = 6.45×10-14 M


Answer: [H3O+] = 0.155 M; [OH-] = 6.45×10-14 M.